# On Exact Space-Depth Trade-Offs in Multi-Controlled Toffoli Decomposition

### v2

@@ Abstract
In this paper, we consider the optimized implementation of Multi Controlled Toffoli (MCT) using the Clifford $+$ T gate sets. While there are several recent works in this direction, here we explicitly quantify the trade-off (with concrete formulae) between the Toffoli depth (this means the depth using the classical 2-controlled Toffoli) of the $n$-controlled Toffoli (hereform we will tell $n$-MCT) and the number of clean ancilla qubits. Additionally, we achieve a reduced Toffoli depth (and consequently, T-depth), which is an extension of the technique introduced by Khattar et al. (2024). In terms of a negative result, we first show that using such conditionally clean ancilla techniques, Toffoli depth can never achieve exactly $\lceil\log_{2}n\rceil$, though it remains of the same order. This highlights the limitation of the techniques exploiting conditionally clean ancilla [Nie et al., 2024, Khattar et al., 2024]. Then we prove that, in a more general setup, the T-Depth in the Clifford + T decomposition, via Toffoli gates, is lower bounded by $\lceil\log_{2}n\rceil$, and this bound is achieved following the complete binary tree structure. Since the ($2$-controlled) Toffoli gate can further be decomposed using Clifford $+$ T, various methodologies are explored too in this regard for trade-off related implications.
@@ I Introduction
Quantum gates are the fundamental building blocks of quantum circuits. Unlike classical gates, quantum gates are inherently reversible and are mathematically represented by unitary matrices. The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations, reversible computing, and oracle constructions wang2024boosting ; wang2024optimal ; wang2025comprehensive ; wang2024Minimum ; wang2024Reducing ; wang2023higher . However, the Toffoli gate’s high resource demands, particularly in terms of T-Count, T-Depth, and ancilla qubits, can greatly influence the efficiency of fault-tolerant quantum circuits. Consequently, optimizing its implementation is crucial for reducing computational overhead, minimizing error rates, and enhancing scalability, thereby making it indispensable for practical large-scale quantum computations.
@@
Over the past decades, numerous efforts have been made to reduce the resource requirements for the optimized implementation of multi-controlled Toffoli (MCT) gates miller11 ; saeedi13 ; maslov16 ; baker19 ; paler22 ; bala22 ; claudon24 . More recent works nie24 ; khattar24 have introduced the conditionally clean ancilla technique, which significantly reduces both the Toffoli depth and ancilla count in the MCT decomposition.
Before proceeding further, let us explain the concept of conditionally clean ancilla that will be repeatedly referred to in this work. Generally, ancilla qubits are of two types, namely the clean ancilla, which is initialized to |$\ket{0}$ at the beginning of the circuit, and the dirty ancilla, which is initialized to some unknown quantum state at the beginning. The standard practice is to re-initialize the ancilla qubit to its initial state at the end of the computation. Conditionally clean ancilla nie24 ; khattar24 is the set of working (control) qubits, derived from the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation. Although earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancilla derived from the clean ones only. Very recently, the conditionally clean ancilla technique has gained significant attention, including its applications in quantum adders remaud25 .
@@
In this paper, we explore how the Toffoli depth (and consequently, T-Depth) can be reduced by increasing the number of clean ancilla qubits, utilizing the concept of conditionally clean ancilla. We also establish the limitation of this method in terms of the lower bound of the Toffoli depth. Additionally, we show that, in a more general setting, the Toffoli depth (which can be further reduced to T-Depth) of an $n$-MCT decomposition cannot be reduced beyond $\lceil\log_{2}n\rceil$. The section-wise contributions of this paper are outlined as follows.
@@ I.1 Organization and Contributions
In section IV, we explore the trade-off between the (clean) ancilla and the Toffoli depth using the existing techniques related to conditionally clean ones. In Section IV.1, we take a different look at viewing the MCT circuit decomposition using the conditionally clean ancilla technique of khattar24 and enumerate their exact Toffoli count. Following the trade-off, we show the reduction in Toffoli depth, and therefore in T-Depth (Construction 1, Example 1) compared to the recent work of Khattar and Gidney khattar24 , by introducing additional clean ancilla into the circuit, while keeping the Toffoli count constant. These results are shown in Section IV.2. Then, in Section IV.3, we identify the limitation of this technique in Theorem 3, showing that this direction cannot reduce the Toffoli depth to $\lceil\log_{2}n\rceil$, though it is of order $\mathcal{O}(\log_{2}n)$.
In Section V, we prove within a general framework that the exact Toffoli depth in the ($2$-controlled) Toffoli decomposition of an $n$-MCT cannot be less than $\lceil\log_{2}n\rceil$, regardless of the number of ancilla qubits used. In fact, this lower bound can be achieved for T-Depth as well, by the construction of jaques19 . That is, using the technique of jaques19 , we can obtain an $n$-MCT by further decomposing into the Clifford + T gate set with an exact T-Depth of $\lceil\log_{2}n\rceil$ too, using $22n-$ 2 ancilla, and a T-Count of $4(n-1)$. Moreover, using the logical-AND circuit by Gidney gidney18 , the ancilla count can be reduced to $n-2$ keeping the T-Count constant. However, the exact T-Depth in the case of gidney18 becomes one more, i.e., $\lceil\log_{2}n\rceil+1$. To highlight our contribution, we are looking at the exact counts and depth instead of their complexity order.
@@ II Preliminaries
In quantum computing, the Multi-Controlled Toffoli (MCT) gates play a crucial role in the design of complex quantum algorithms, including error correction codes and arithmetic operations. Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancillary qubits. These optimizations are essential for practical quantum computation, where resource efficiency is a critical consideration. In this regard, we present a comprehensive overview of the existing benchmarks for Clifford + T decompositions of single and multi-controlled Toffoli gates in the following section.
Let us now describe the idea of conditionally clean ancilla from khattar24 , as described in Figure 2. Given a clean ancilla, a Toffoli gate is implemented targeting the clean ancilla. If the clean ancilla is reversed, it means both the control qubits were |$\ket{1}$; thus, applying X gates on the control qubits will change them to |$\ket{0}$ and consequently can be used as conditionally clean ancilla in the next rounds. Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancilla, and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits. Once the information from all the control qubits is accumulated in a few qubits, another (smaller) MCT gate is implemented with these qubits along with the ancilla as the control qubits and the original target as its target, so that if the ancilla was not modified in the first step, the target will also not be modified.
Figure 2: Understanding the conditionally clean ancilla technique from khattar24 .
@@
Case II: At least one of $C_{i}=\ket{0}$, for some $i\in[0,4]$. If one of $C_{0}$ or $C_{1}$ is |$\ket{0}$, then ancilla =|$\ket{0}$, and the target will not be flipped in Step III. If $C_{0}=C_{1}=\ket{1}$, but one of $C_{2},C_{3},C_{4}$ is |$\ket{0}$, then ancilla becomes |$\ket{1}$, and after applying X-gate on $C_{0},C_{1}$ will make, $C_{0}=C_{1}=\ket{0}$. If one of $C_{2},C_{3}$ is |$\ket{0}$, then $C_{0}$ and consequently, $C_{1}$ also remains at |$\ket{0}$ state. Similarly, if $C_{4}=\ket{0}$, then also $C_{1}$ remains |$\ket{0}$, and in both scenario, the target will not be flipped in Step III.
In Section IV.2, we modify the design by introducing additional ancilla qubits in Step I and subsequently creating more conditionally clean ancilla to begin with, which eventually reduces the overall Toffoli depth of the complete circuit.
@@ III.1 Single Toffoli Decomposition
To immediately dive into the technical issues, as shown in (nc, , Ch. 4, Sec. 4.3), there exists a Clifford + T decomposition of a single Toffoli gate using $7$ T gates and $9$ Clifford gates (2 H, 1 S, 6 CNOT), with a T-Depth of $6$ (Figure 3(a)). Through a simple manipulation of gate ordering, the T-Depth can be reduced to $4$ without altering the gate count. In 2013, Amy et al. (amy13, , Section 6) proposed the Toffoli decomposition (Figure 3(b)) with a T-Depth of $4$ that used one fewer Clifford gate and reduced the overall circuit depth from $12$ to $8$. In the same paper, the authors applied a meet-in-the-middle algorithm to present a Toffoli decomposition (Figure 3(c)) with a T-Depth of $3$ and an overall depth of $9$. For an exact decomposition of a single Toffoli gate (without measurement-based feedback), this is the lowest T-Depth that one can achieve without using any ancillary qubit.
@@ III.2 Multi-Controlled Toffoli Decomposition
Figure 10: MCT decomposition using conditionally clean ancilla due to nie24 .
@@
Furthermore, when the availability of clean ancilla increases to $2$ (Figure 11(b)), the Toffoli depth reduces to $\mathcal{O}(\log_{2}n)$ while maintaining the Toffoli count constant. Figure 11 illustrates the circuit diagrams of $10$-MCT using $1$ and $2$ clean ancilla, due to khattar24 . Additionally, if the clean ancilla is replaced with the dirty ancilla, the Toffoli count increases to $3216n-$ 32, and the Toffoli depth doubles under both scenarios. As we focus here on the conditionally clean ancilla derived solely from clean ancilla, we do not delve into an exact analysis of the Toffoli depth for the dirty ancilla circuit implementation.
@@
In this section, we provided the exact enumerations of Toffoli count and Toffoli depth of MCT circuit decomposition across various state-of-the-art results. For the exact enumeration of the Toffoli depth (and T-Depth), following khattar24 , we propose a novel approach of viewing the MCT decomposition with conditionally clean ancilla, which is a direct extension of khattar24 , thereby presented separately in Proposition 3 of Section IV.1. Later, the approach has also been used to demonstrate the trade-off between Toffoli depth and the clean ancilla count.
@@ IV Further Reduction of Toffoli Depth
In this section, we adopt the MCT decomposition technique using the conditionally clean ancilla, as introduced in nie24 ; khattar24 , and propose a trade-off between the Toffoli depth with the clean ancilla, showing improvement over the recent work of Khattar and Gidney khattar24 . The authors proposed the MCT circuit decompositions techniques, achieving a Toffoli depth of $\mathcal{O}(n)$ with a single clean ancilla and $\mathcal{O}(\log_{2}n)$ with two clean ancillae, where in both cases, the Toffoli count becomes $32n-$ 3. Here, we revisit the MCT decompositions by khattar24 and provide an exact enumeration of the Toffoli depth (consequently T-Depth) that was not explicitly presented in khattar24 .
@@ IV.1 Exact Enumeration of Toffoli Depth proposed by Khattar and Gidney (2024) in a Different Lens
The $n$-MCT decomposition circuit khattar24 using two ancilla qubits consists of five steps. Step I and Step V combined have a Toffoli depth of 1 (using logical-AND). In Step II, information from the remaining $n-2$ control qubits is accumulated into $k(\ll n)$ qubits using the conditionally cleaned ancilla technique. Next, in Step III, a $(k+1)$-MCT is implemented using the unused ancilla, following the single-ancilla technique, having a Toffoli depth of $12(k+1)-3=2k-$ 1. In Step IV, the control qubits are returned to their original state through uncomputation, thereby having the same Toffoli count and depth as in Step II.
@@
In the above example, Step I and Step II of a $32$-MCT decomposition have been shown using the conditionally cleaned ancilla technique. The Step I has a Toffoli depth of $1$, and Step II has a Toffoli depth of $7$. Further, in Step III, a $5$-MCT needs to be applied to hit the target qubit, having a Toffoli depth $2(5)-3=7$. However, a part of the smaller MCT is applied simultaneously with Step II and Step IV, thereby having an effective Toffoli depth of $3$ or $5$. Clearly, the uncomputation (Step IV) requires an additional Toffoli depth of $8$, i.e., the Toffoli depth of the complete circuit is $19$.
Using this new approach, the $5$-MCT circuit using a conditionally clean ancilla technique, presented in Figure 2, can be seen as follows.
@@
In the direction of enumerating the exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancilla, we have the following lemma.
@@
Following khattar24 , the exact Toffoli depth of an $n$-MCT decomposition using $2$ clean ancilla is lower bounded by $3\log_{2}n$.
@@
Although the above proposition establishes that the exact Toffoli depth using two clean ancillae, as per khattar24 , is lower bounded by $3\log_{2}n$, this bound is not tight and is closer to $4\log_{2}n$. In the following section, we analyze the trade-off between Toffoli depth and clean ancilla, demonstrating that with additional ancilla, the exact Toffoli depth can be reduced to $2\log_{2}n$.

### v3

@@ Abstract
In this paper, we consider the optimized implementation of Multi Controlled Toffoli (MCT) using the Clifford $+$ T gate sets. While there are several recent works in this direction, here we explicitly quantify the trade-off (with concrete formulae) between the Toffoli depth (this means the depth using the classical 2-controlled Toffoli) of the $n$-controlled Toffoli (hereform we will tell $n$-MCT) and the number of clean ancilla qubits. Additionally, we achieve a reduced Toffoli depth (and consequently, T-depth), which is an extension of the technique introduced by Khattar et al. (2024). In terms of a negative result, we first show that using such conditionally clean ancillae techniques, Toffoli depth can never achieve exactly $\lceil\log_{2}n\rceil$, though it remains of the same order. This highlights the limitation of the techniques exploiting conditionally clean ancillae [Nie et al., 2024, Khattar et al., 2024]. Then we prove that, in a more general setup, the T-Depth in the Clifford + T decomposition, via Toffoli gates, is lower bounded by $\lceil\log_{2}n\rceil$, and this bound is achieved following the complete binary tree structure. Since the ($2$-controlled) Toffoli gate can further be decomposed using Clifford $+$ T, various methodologies are explored too in this regard for trade-off related implications.
@@ I Introduction
Quantum gates are the fundamental building blocks of quantum circuits. Unlike classical gates, quantum gates are inherently reversible and are mathematically represented by unitary matrices. The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations wang2024boosting ; wang2024optimal ; wang2024Minimum ; wang2023higher , reversible computing lighter2019 ; dorcis2023 , and oracle constructions aes16 ; chacha21 ; fsr21 ; zuc24 . However, the Toffoli gate’s high resource demands, particularly in terms of T-Count, T-Depth, and ancilla qubits, can greatly influence the efficiency of fault-tolerant quantum circuits. Consequently, optimizing its implementation is crucial for reducing computational overhead, minimizing error rates, and enhancing scalability, thereby making it indispensable for practical large-scale quantum computations.
@@
Over the past decades, numerous efforts have been made to reduce the resource requirements for the optimized implementation of multi-controlled Toffoli (MCT) gates miller11 ; saeedi13 ; maslov16 ; baker19 ; paler22 ; bala22 ; claudon24 . More recent works nie24 ; khattar24 have introduced the conditionally clean ancillae technique, which significantly reduces both the Toffoli depth and ancilla count in the MCT decomposition.
Before proceeding further, let us explain the concept of conditionally clean ancillae that will be repeatedly referred to in this work. Generally, ancilla qubits are of two types, namely the clean ancilla, which is initialized to |$\ket{0}$ at the beginning of the circuit, and the dirty ancilla, which is initialized to some unknown quantum state at the beginning. The standard practice is to re-initialize the ancilla qubit to its initial state at the end of the computation. Conditionally clean ancillae nie24 ; khattar24 are the set of working (control) qubits, derived from the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation. Although earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancillae derived from the clean ones only. Very recently, the conditionally clean ancillae technique has gained significant attention, including its applications in quantum adders remaud25 .
@@
In this paper, we explore how the Toffoli depth (and consequently, T-Depth) can be reduced by increasing the number of clean ancilla qubits, utilizing the concept of conditionally clean ancillae. We also establish the limitation of this method in terms of the lower bound of the Toffoli depth. Additionally, we show that, in a more general setting, the Toffoli depth (which can be further reduced to T-Depth) of an $n$-MCT decomposition cannot be reduced beyond $\lceil\log_{2}n\rceil$. The section-wise contributions of this paper are outlined as follows.
@@ I.1 Organization and Contributions
In section IV, we explore the trade-off between the (clean) ancilla and the Toffoli depth using the existing techniques related to conditionally clean ones. In Section IV.1, we take a different look at viewing the MCT circuit decomposition using the conditionally clean ancillae technique of khattar24 and enumerate their exact Toffoli count. Following the trade-off, we show the reduction in Toffoli depth, and therefore in T-Depth (Construction 1, Example 1) compared to the recent work of Khattar and Gidney khattar24 , by introducing additional clean ancillae into the circuit, while keeping the Toffoli count constant. These results are shown in Section IV.2. Then, in Section IV.3, we identify the limitation of this technique in Theorem 3, showing that this direction cannot reduce the Toffoli depth to $\lceil\log_{2}n\rceil$, though it is of order $\mathcal{O}(\log_{2}n)$.
In Section V, we prove within a general framework that the exact Toffoli depth in the ($2$-controlled) Toffoli decomposition of an $n$-MCT cannot be less than $\lceil\log_{2}n\rceil$, regardless of the number of ancilla qubits used. In fact, this lower bound can be achieved for T-Depth as well, by the construction of jaques19 . That is, using the technique of jaques19 , we can obtain an $n$-MCT by further decomposing into the Clifford + T gate set with an exact T-Depth of $\lceil\log_{2}n\rceil$ too, using $22n-$ 2 ancillae, and a T-Count of $4(n-1)$. Moreover, using the logical-AND circuit by Gidney gidney18 , the ancilla count can be reduced to $n-2$ keeping the T-Count constant. However, the exact T-Depth in the case of gidney18 becomes one more, i.e., $\lceil\log_{2}n\rceil+1$. To highlight our contribution, we are looking at the exact counts and depth instead of their complexity order.
@@ II Preliminaries
In quantum computing, the Multi-Controlled Toffoli (MCT) gates play a crucial role in the design of complex quantum algorithms, including error correction codes and arithmetic operations. Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancilla qubits. These optimizations are essential for practical quantum computation, where resource efficiency is a critical consideration. In this regard, we present a comprehensive overview of the existing benchmarks for Clifford + T decompositions of single and multi-controlled Toffoli gates in the following section.
Let us now describe the idea of conditionally clean ancillae from khattar24 , as described in Figure 2. Given a clean ancilla, a Toffoli gate is implemented targeting the clean ancilla. If the clean ancilla is reversed, it means both the control qubits were |$\ket{1}$; thus, applying X gates on the control qubits will change them to |$\ket{0}$ and consequently can be used as conditionally clean ancillae in the next rounds. Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancillae, and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits. Once the information from all the control qubits is accumulated in a few qubits, another (smaller) MCT gate is implemented with these qubits along with the ancilla as the control qubits and the original target as its target, so that if the ancilla was not modified in the first step, the target will also not be modified.
Figure 2: Understanding the conditionally clean ancillae technique from khattar24 .
@@
Case II: At least one of $C_{i}=\ket{0}$, for some $i\in[0,4]$. If one of $C_{0}$ or $C_{1}$ is |$\ket{0}$, then ancilla =|$\ket{0}$, and the target will not be flipped in Step III. If $C_{0}=C_{1}=\ket{1}$, but one of $C_{2},C_{3},C_{4}$ is |$\ket{0}$, then the ancilla becomes |$\ket{1}$, and after applying X-gate on $C_{0},C_{1}$ will make, $C_{0}=C_{1}=\ket{0}$. If one of $C_{2},C_{3}$ is |$\ket{0}$, then $C_{0}$ and consequently, $C_{1}$ also remains at |$\ket{0}$ state. Similarly, if $C_{4}=\ket{0}$, then also $C_{1}$ remains |$\ket{0}$, and in both scenario, the target will not be flipped in Step III.
In Section IV.2, we modify the design by introducing additional ancilla qubits in Step I and subsequently creating more conditionally clean ancillae to begin with, which eventually reduces the overall Toffoli depth of the complete circuit.
@@ III.1 Single Toffoli Decomposition
To immediately dive into the technical issues, as shown in (nc, , Ch. 4, Sec. 4.3), there exists a Clifford + T decomposition of a single Toffoli gate using $7$ T gates and $9$ Clifford gates (2 H, 1 S, 6 CNOT), with a T-Depth of $6$ (Figure 3(a)). Through a simple manipulation of gate ordering, the T-Depth can be reduced to $4$ without altering the gate count. In 2013, Amy et al. (amy13, , Section 6) proposed the Toffoli decomposition (Figure 3(b)) with a T-Depth of $4$ that used one fewer Clifford gate and reduced the overall circuit depth from $12$ to $8$. In the same paper, the authors applied a meet-in-the-middle algorithm to present a Toffoli decomposition (Figure 3(c)) with a T-Depth of $3$ and an overall depth of $9$. For an exact decomposition of a single Toffoli gate (without measurement-based feedback), this is the lowest T-Depth that one can achieve without using any ancilla qubit.
@@ III.2 Multi-Controlled Toffoli Decomposition
Figure 10: MCT decomposition using conditionally clean ancillae due to nie24 .
@@
Furthermore, when the availability of clean ancilla qubit increases to $2$ (Figure 11(b)), the Toffoli depth reduces to $\mathcal{O}(\log_{2}n)$ while maintaining the Toffoli count constant. Figure 11 illustrates the circuit diagrams of $10$-MCT using $1$ and $2$ clean ancillae, due to khattar24 . Additionally, if the clean ancillae is replaced with the dirty ancillae, the Toffoli count increases to $3216n-$ 32, and the Toffoli depth doubles under both scenarios. As we focus here on the conditionally clean ancillae derived solely from clean ancilla qubits, we do not delve into an exact analysis of the Toffoli depth for the dirty ancilla circuit implementation.
@@
In this section, we provided the exact enumerations of Toffoli count and Toffoli depth of MCT circuit decomposition across various state-of-the-art results. For the exact enumeration of the Toffoli depth (and T-Depth), following khattar24 , we propose a novel approach of viewing the MCT decomposition with conditionally clean ancillae, which is a direct extension of khattar24 , thereby presented separately in Proposition 3 of Section IV.1. Later, the approach has also been used to demonstrate the trade-off between Toffoli depth and the clean ancilla count.
@@ IV Further Reduction of Toffoli Depth
In this section, we adopt the MCT decomposition technique using the conditionally clean ancillae, as introduced in nie24 ; khattar24 , and propose a trade-off between the Toffoli depth with the clean ancilla, showing improvement over the recent work of Khattar and Gidney khattar24 . The authors proposed the MCT circuit decompositions techniques, achieving a Toffoli depth of $\mathcal{O}(n)$ with a single clean ancilla and $\mathcal{O}(\log_{2}n)$ with two clean ancillae, where in both cases, the Toffoli count becomes $32n-$ 3. Here, we revisit the MCT decompositions by khattar24 and provide an exact enumeration of the Toffoli depth (consequently T-Depth) that was not explicitly presented in khattar24 .
@@ IV.1 Exact Enumeration of Toffoli Depth proposed by Khattar and Gidney (2024) in a Different Lens
The $n$-MCT decomposition circuit khattar24 using two ancilla qubits consists of five steps. Step I and Step V combined have a Toffoli depth of 1 (using logical-AND). In Step II, information from the remaining $n-2$ control qubits is accumulated into $k(\ll n)$ qubits using the conditionally cleaned ancillae technique. Next, in Step III, a $(k+1)$-MCT is implemented using the unused ancilla, following the single-ancilla technique, having a Toffoli depth of $12(k+1)-3=2k-$ 1. In Step IV, the control qubits are returned to their original state through uncomputation, thereby having the same Toffoli count and depth as in Step II.
@@
In the above example, Step I and Step II of a $32$-MCT decomposition have been shown using the conditionally cleaned ancillae technique. The Step I has a Toffoli depth of $1$, and Step II has a Toffoli depth of $7$. Further, in Step III, a $5$-MCT needs to be applied to hit the target qubit, having a Toffoli depth $2(5)-3=7$. However, a part of the smaller MCT is applied simultaneously with Step II and Step IV, thereby having an effective Toffoli depth of $3$ or $5$. Clearly, the uncomputation (Step IV) requires an additional Toffoli depth of $8$, i.e., the Toffoli depth of the complete circuit is $19$.
Using this new approach, the $5$-MCT circuit using a conditionally clean ancillae technique, presented in Figure 2, can be seen as follows.
@@
In the direction of enumerating the exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancillae, we have the following lemma.
@@
Following khattar24 , the exact Toffoli depth of an $n$-MCT decomposition using $2$ clean ancillae is lower bounded by $3\log_{2}n$.
@@
Although the above proposition establishes that the exact Toffoli depth using two clean ancillae, as per khattar24 , is lower bounded by $3\log_{2}n$, this bound is not tight and is closer to $4\log_{2}n$. In the following section, we analyze the trade-off between Toffoli depth and clean ancilla, demonstrating that with additional ancilla qubit, the exact Toffoli depth can be reduced to $2\log_{2}n$.